40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 4 20:13:39 2019, gautam, Update, SUS, ETMY suspension oddness ETMY_sensors_1_Feb_2019_2230_PST.pdfETMY_UL.pdfETMY_sensors_timeDomain.pdf
    Reply  Thu Feb 7 19:34:18 2019, gautam, Update, SUS, ETMY suspension oddness ETMY_sensors.pdfpostDiag.pdf
       Reply  Fri Feb 8 20:48:52 2019, gautam, Update, LSC, IFO recovery IMG_7307.JPG
          Reply  Mon Feb 11 15:41:49 2019, gautam, Update, LSC, TRY 60 Hz solved - but clipping persists 
             Reply  Mon Feb 11 16:38:34 2019, gautam, Update, LSC, ETMY OL calibration updated OL_calib_ETMY_PERROR.pdfOL_calib_ETMY_YERROR.pdf
             Reply  Fri Feb 15 18:42:57 2019, gautam, Update, LSC, TRY 60 Hz solved  IMG_7323.JPG
Message ID: 14433     Entry time: Mon Feb 4 20:13:39 2019     Reply to this: 14441
Author: gautam 
Type: Update 
Category: SUS 
Subject: ETMY suspension oddness 

I looked at the free-swinging sensor data from two nights ago, and am struggling with the interpretation. 

[Attachment #1] - Fine resolution spectral densities of the 5 shadow sensor signals (y-axis assumes 1ct ~1um). The puzzling feature is that there are only 3 resonant peaks visible around the 1 Hz region, whereas we would expect 4 (PIT, YAW, POS and SIDE). afaik, Lydia looked into the ETMY suspension diagonalization last, in 2016. Compared to her plots (which are in the Euler basis while mine are in the OSEM basis), the ~0.73 Hz peak is nowhere to be seen. I also think the frequency resolution (<1 mHz) is good enough to be able to resolve two closely spaced peaks, so it looks like due to some reason (mechanical or otherwise), there are only 3 independent modes being sensed around 1 Hz.

[Attachment #2] - Koji arrived and we looked at some transfer functions to see if we could make sense of all this. During this investigation, we also think that the UL coil actuator electronics chain has some problem. This test was done by driving the individual coils and looking for the 1/f^2 pendulum transfer function shape in the Oplev error signals. The ~ 4dB difference between UR/LL and LR is due to a gain imbalance in the coil output filter bank, once we have solved the other problems, we can reset the individual coil balancing using this measurement technique.

[Attachment #3] - Downsampled time-series of the data used to make Attachment #1. The ringdown looks pretty clean, I don't see any evidence of any stuck magnets looking at these signals. The X-axis is in kilo-seconds.

We found that the POS and SIDE local damping loops do not result in instability building up. So one option is to use only Oplevs for angular control, while using shadow-sensor damping for POS and SIDE.

Attachment 1: ETMY_sensors_1_Feb_2019_2230_PST.pdf  124 kB  | Hide | Hide all
ETMY_sensors_1_Feb_2019_2230_PST.pdf
Attachment 2: ETMY_UL.pdf  7 kB  | Hide | Hide all
ETMY_UL.pdf
Attachment 3: ETMY_sensors_timeDomain.pdf  40 kB  Uploaded Mon Feb 4 23:03:23 2019  | Hide | Hide all
ETMY_sensors_timeDomain.pdf
ELOG V3.1.3-